# Kiloparsec-scale turbulence driven by reionization may grow intergalactic magnetic fields

### v1

@@ Abstract
The intergalactic medium (IGM) underwent intense heating that resulted in pressure disequilibrium in the wake of ionization fronts during cosmic reionization. The dynamical relaxation to restore pressure balance may have driven small-scale turbulence and, hence, the amplification of intergalactic magnetic fields. We investigate this possibility using a suite of $\approx 100$ pc resolution radiation-hydrodynamics simulations of IGM gas dynamics. We show that as the spatial resolution improves beyond that achieved with most prior studies, much of the IGM becomes turbulent unless it was pre-heated to $\gg 100~{}$K before reionization. In our most turbulent simulations, we find that the gas energy spectrum follows the expected $k^{-5/3}$ Kolmogorov scaling to the simulation’s resolution, and the eddy turnover time of the turbulence is $<1$ Gyr at $k\approx 1~{}$kpc-1. Turbulence will grow magnetic fields, and we show that the fields grown by reionization-driven turbulence could explain lower limits on IGM B-field strengths from observations of TeV blazars.
@@ I Introduction
Hydrogen in the Intergalactic Medium (IGM) underwent a major phase transition about a billion years after the big bang, when it was ionized by the first generation of galaxies and possibly active galactic nuclei. This transition not only changed its ionization state, but also its temperature [1, 2, 3] and small-scale dynamics [4, 5, 6, 7]. Prior to becoming reionized, models predict that the IGM temperature was in the range $10$-100010001000~{}1000[8, 9, 10]. For the coldest of these temperatures, the gas may have formed structures as light as $10^{4}$ $M_{\odot}$ [11]. Reionization impulsively heated this gas to a few ×$\times 10^{4}$, over-pressurizing these systems and causing them to dynamically relax to a scale of tens of kiloparsecs over a few hundred million years [5].
@@
Several studies have used high-resolution cosmological simulations to study this relaxation process and how it affects various observables  [12, 4, 5, 13, 14, 15, 16]. Here we investigate whether relaxation could drive turbulence that was unresolved in these prior studies. There is reason to suspect that the post-reionization IGM may have been turbulent. After ionization fronts (I-fronts) sweep through a region, the gas would have been evacuated out of mini-halos and filaments and transported $\sim 10$s of co-moving kiloparsecs [5, 7], providing a natural driving mechanism. As the recently ionized gas expands out of the potential well, it cools adiabatically, reaching temperatures as low as a few thousand Kelvin [13, 14]. Nearby diffuse gas is compressed by the expanding over-densities, and can reach $5\times 10^{4}~{}$ or even hotter within $\approx 100$ Myr of reionization. The pressure differences between hot and cool gas can then drive turbulence.
@@
Turbulent conditions in the IGM during and after reionization have the potential to amplify intergalactic magnetic fields. There has been significant previous research on the seeding of IGM magnetic fields. One of the most referenced seeding mechanisms is during reionization itself: the inhomogeneous heating leads to small charge segregation and the resulting electric fields generate fields via Faraday’s law, the “Biermann battery” [17, 18, 19]. Additionally, there may even be seeds generated prior to reionization, such as from the streaming of cosmic rays produced by the first supernovas [20] and possibly even primordially during the Big Bang [21]. However, with the exception of primordial fields which generally require novel fundamental particles, known seeding mechanisms predict seed fields of $\lesssim 10^{-20}$ G. Reaching larger field strengths requires an amplification mechanism. One such mechanism is the so-called turbulent dynamo, which has been explored in detail both in general setups [22, 23, 24] and, for example, in the context of forming proto-galaxies [25, 26, 23]. Here we consider the turbulent dynamo in the context of reionization-driven turbulence.
@@
Explaining the growth of volume-filling IGM magnetic fields is of particular interest in light of recent lower limits on the strength of these fields from TeV Blazars [27, 28]. In the absence of magnetic fields, one expects to observe $\gamma$-ray halos around these objects generated by the inverse-Compton scattering of TeV photons off of CMB photons. The absence of these halos both observed in the spectral energy distribution and the spatial extent of the sources has been interpreted as an effect of deflection by IGM magnetic fields (an alternative plasma instability hypothesis has been recently ruled out  [^1] An alternative explanation of the blazar observations is that the TeV pairs lose their energy owing to plasma beam instabilities [55]. After an extensive debate, this possibility appears to have been ruled out. It was recently shown that Landau damping by non-thermal electrons generated by Compton scattering off of the gamma ray background should suppress the growth of these instabilities [56].). This absence places a conservative lower limit on their field strength of $B>3\times 10^{-16}$ G for fields correlated over $10$ kpc [30, 31, 32]. A non-primordial explanation for these constraints requires significant amplification of seed fields throughout the IGM. We explore in this Letter whether this can be acheived by turbulence driven by reionization  [^2] Galactic and AGN feedback might be the other possibility [57], although this is not anticipated to be volume filling..
@@ II Numerical Methods
Our simulations of small-scale IGM gas dynamics were run using a modified version of the RadHydro code of Trac and Pen [35] and Trac and Cen [36]. Our version of RadHydro performs plane-parallel ray tracing radiative transfer (RT), with rays traced from the boundaries of cubical sub-volumes (RT domains) within our simulation box. Gravity is solved using the particle-mesh (PM) method for computational efficiency. Our simulations are initialized at $z=300$ using separate transfer functions for dark matter and baryons from CAMB [37] and are run to $z=4$. Our fiducial box size is $L_{\rm box}=250$ $h^{-1}$kpc, with $1024^{3}$ gas cells, RT cells, and dark matter particles, for a fiducial spatial resolution of $\Delta x=244$ $h^{-1}$pc. Plane-parallel ionization fronts (I-fonts) cross the RT domains at a redshift $z_{\rm re}$, following the procedure described in Cain et al. [14] (see also Appendix B). We use the reduced speed of light approximation to save computation time after I-fronts have crossed the RT domains (see below). Following D’Aloisio et al. [5], the ionizing spectrum is discretized into $5$ frequency bins between $1$ and $4$ Ry, with a power law frequency dependence $J_{\nu}\propto\nu^{-1.5}$. In our fiducial simulations, the strength of the ionizing background is set such that the HI photo-ionization rate is $\Gamma_{\rm HI}=3\times 10^{-13}$ s-1 everywhere except within self-shielded systems. To study numerical convergence, we have also run a simulation with $2048^{3}$ grid cells and dark matter particles ($\Delta x=122$ $h^{-1}$pc), and ones with lower resolution. Due to computational limitations, we flash-ionized this $2048^{3}$ run instead of tracking RT. We show in Appendix B that this approximation does not significantly alter the subsequent relaxation.
Figure 1: Time series of snapshots showing the evolution of the density (top panels) and temperature (bottom panels) in our fiducial $0.25~{}h^{-1}~{}$pc, $1024^{3}$ simulation. Immediately after ionization ($z=6.99$), the gas is clumpy down to sub-kiloparsec scales, reflecting its pre-ionized cold temperature, and is nearly isothermal, reflecting the heat input from reionization. Subsequently, over-dense filaments expand, generating a medium of expansion-cooled and compression-heated gas. By $z=6$, this dynamics beings to result in turbulence, and at $z\leq 5.5$, turbulence is observed throughout, in both the gas density and temperature. As the dynamical relaxation slows, the turbulence dissipates, but some still persists to at least to $z=4$.
@@ III.1 Dynamics
Figure 1 shows the time evolution of the IGM gas density (top panels) and temperature (bottom panels) in our fiducial simulation, where reionization occurs at a redshift of $z_{\rm reion}=7$. At $z=6.99$, immediately after ionization, the gas has had negligible time to respond to the photo-heating and remains clumpy down to sub-kpc scales. The low density gas is nearly isothermal, at $\approx 30,000~{}$K, while the over-dense filaments are slightly cooler. These temperature differences result from the I-fronts moving more slowly across the filaments, since the resulting temperature scales with the I-front speed  [2, 3].
@@
Figure 2: Comparison of turbulence in runs with and without X-ray pre-heating, showing our highest resolution $0.25~{}h^{-1}~{}$pc, $2048^{3}$ simulations. Left: slice through the temperature field at $z=5$. The simulation assumes no heating before reionization such that the evolution is essentially adiabatic. Right: the same, but for a run with a minimum gas temperature of $100$ K imposed at $z<15$. We see less turbulence in this case, with it being concentrated largely around the filaments.
@@
Between $z=7$ and $6.5$, the gas evolves dynamically in a way consistent with the findings of Cain et al. [14] (see also Hirata [13]). The filamentary structures expand and cool adiabatically in response to impulsive heating from reionization. Meanwhile, the gas surrounding the expanding filaments is compressed and heated to $\gtrsim 50,000~{}$K. By $z=6.5$, there is a factor of $\approx 10$ temperature difference between over-dense, expanding gas and surrounding hot, compressed gas at lower densities. At $z=6$, these two phases are beginning to turbulently mix, and, by $z=5.5$, turbulence has developed, leading to significant small-scale structure in both the gas density and temperature. Some of the turbulence may be instigated by Rayleigh-Taylor instabilities of less dense but over-pressurized gas expanding into denser but lower pressure regions. By $z=4$, the pressure relaxation is nearly complete, and the turbulence has started to dissipate. Still, turbulent structure is evident down to this redshift, $\approx 700$ Myr after the gas was reionized.
@@
No previous works (to our knowledge) have observed turbulent behavior in the context of the reionized IGM, absent the turbulence driven by Galactic winds near galaxies. We believe this is because they did not achieve the required spatial resolution. The highest-resolution studies to-date in this context (using grid or particle-based codes) achieve resolutions of $\approx$ $1-2$ ckpc [4, 13, 5, 7]. The focus of these studies was resolving the pre-ionization Jeans scale – the smallest scale at which the gas clumps –, which corresponds to $\approx 100$kpc (see Appendix A of [5]). We have run a convergence test (see supplemental material) and found that turbulent features are not seen in our simulations at all for cell sizes larger than $0.5$ $h^{-1}$kpc. We show in Appendix A that this resolution requirement is set not by the Jeans scale, but by the scale at which numerical discretization introduces effective viscosity that prevents turbulence from forming [38]. We further show that in the context of post-reionization pressure smoothing, resolutions of $0.5-1$ kpc are the expected minimum requirement to resolve turbulence.
@@ III.2 Effect of X-ray pre-heating
Pre-heating by X-rays before reionization increases the Jeans scale of the pre-ionized gas above that of the limit of pure “adiabatic” evolution that is assumed by our fiducial simulation  [^3] Note that Compton heating off the CMB is significant at $z\gtrsim 100$, and we do include this effect. At $z\lesssim 100$, cosmic expansion and X-ray pre-heating at the only significant heating/cooling channels before reionization. . Smoothing of small-scale structure by pre-heating could inhibit the formation of turbulence by eliminating the smallest structures and making it less likely for expanding filaments to overlap. Indeed, upper limits on the 21 cm power spectrum from HERA [40] already constrain the temperature of the neutral IGM to be $\gtrsim 10$ K by $z\approx 8$. Theoretical expectations for the minimum temperature, $T_{\min}$, of the pre-heated IGM are as high as $T_{\min}\approx 1000$[e.g. 41]. However, recent work presenting more careful modeling of the spectra of high-redshift X-ray binaries (the most plausible source of such heating) anticipates temperatures closer to $\approx 100$ K near the start of reionization [8]. We model the effects of X-ray heating in an approximate manner, following D’Aloisio et al. [5], by imposing a temperature floor of $T_{\min}$ at $z<15$. We have run fiducial-resolution simulations with $T_{\min}=10$, $100$, and 100010001000~{}1000K, spanning the range of observational limits and theoretical predictions described above, and one higher-resolution run with $T_{\min}=100~{}$K. We note that our pre-heating implementation likely overestimates the effect of pre-heating on structure formation for a given $T_{\min}$, since X-ray pre-heating is likely to trace the exponential evolution of structure formation [see e.g. 42].
@@
In Figure 2, we show slices through the temperature map at $z=5$ for our highest-resolution (flash-ionized) adiabatic simulation with $122$ $h^{-1}$pc resolution (left), and a run with the same resolution but with $T_{\min}=100$ K (right). As in Figure 1, we see turbulence throughout most of the box in the left panel, particularly surrounding the largest expanding filaments. In the right, however, turbulent features are much more sparse, favoring regions around the most massive filaments. We note that the formation of massive structures, around which turbulence can still form in the $T_{\min}=100$K run, is significantly suppressed in our very small box. As such, the fraction of gas that is turbulent in our $T_{\min}=100~{}$K run is likely an underestimate for the entire IGM. Indeed, the fact that some turbulence is seen in this case given these limitations is an encouraging indicator that it is likely an important effect in the real IGM.
We have run additional tests with $T_{\min}=10$ and $1000$ K (see supplementary material for images) and find that the former is similar to the adiabatic case, while the latter displays no discernible turbulent features. The first result indicates that current lower limits on IGM pre-heating from HERA [40] do not (yet) rule out the level of turbulence seen in our adiabatic simulations. The second suggests that pre-heating to temperatures on the high end of theoretical expectations may be enough to substantially reduce or even eliminate reionization-driven turbulence in the diffuse IGM. We emphasize again that our simplistic treatment of pre-heating and our tiny simulation volumes make this a conservative result. It is also unclear how going to still-higher resolutions or using different hydrodynamics codes (including Eulerian and SPH) that are less numerically diffusive (see Appendix C) would affect these findings.
IV Implications for Intergalactic Magnetic Fields
IV.1 Quantifying turbulence
Let us first quantify the properties of turbulence in our simulations. The turnover timescale at wavenumber $k=2\pi/L$ in a turbulent flow is
$\tau_{\rm eddy}(k)\approx\frac{L}{\sigma_{\rm v}(k)}=\frac{2\pi/k}{[2kE(k^{\prime})]^{1/2}},$ (1)
$E(k)\equiv\sum_{i}\Delta_{\rm v_{i}}^{2}(k)/(2k)$ (2)
is the so-called “energy spectrum”, where $\Delta_{\rm v_{i}}^{2}$ is the dimensionless power spectrum of the $i^{\rm th}$ velocity component. Many studies have shown that the Kolmogorov scaling of $E(k)\propto k^{-5/3}$ tends to be followed over the turbulent inertial range for isotropic, homogeneous turbulence [e.g. 43]. The total number of turnover times elapsed at scale $k$, $N_{\rm\tau}$, is given by integrating $\tau_{\rm eddy}^{-1}$ over time, starting from the time a sound wave can cross the distance scale $2\pi/k$ since the start of reionization.
Figure 3: Quantification of turbulence in our fiducial simulation. Top: The energy spectrum $E(k)$ at a time-series of redshifts, offset vertically for visualization, with redshift decreasing from bottom to top. The “bump” that moves to the left with time tracks the growing sound crossing distance since $z_{\rm re}$. At smaller scales, we find the $k^{-5/3}$ scaling characteristic of turbulence, with a fall-off at very small scales due to numerical dissipation. Bottom: the same as the top panel except without rescaling the amplitude and concentrating on $z=5$ and different spatial resolutions. The high-$k$ fall off shifts to the right as resolution increases, as expected if this cutoff is numerical. The scaling at smaller $k$ is close to, but slightly shallower than, the $-5/3$ Kolmogorov scaling.
In the top left panel of Figure 3, we show $E(k)$ at a number of redshifts (decreasing from bottom to top) for our fiducial simulation. We offset the curves vertically so they can be easily compared, and show the expected $k^{-5/3}$ Kolmogorov scaling as a thin line. We see that just after ionization at $z=6.9$, $E(k)$ has a prominent “bump” centered at $\approx 800$ $h~{}$Mpc-1. The right edge of the bump at $\approx 1500$ $h$ Mpc-1 corresponds to a co-moving distance scale of $\approx 4$ $h^{-1}$kpc, roughly twice the distance that a $20$ km s-1 sound wave can travel through a $T\approx 2\times 10^{4}$ highly photo-ionized gas in the $15$ Myr since $z_{\rm reion}=7$. It is thus reasonable to interpret the bump as tracking the sound crossing distance since reionization. At successively lower redshifts, the bump shifts to smaller $k$ as sound waves travel further from their source, disappearing by $z\approx 5$ as the sound crossing distance approaches the box scale.
We expect turbulence to form at scales smaller than the sound-crossing distance – that is, to the right of the bump. This is exactly what we see – starting at $z=6.5$, $E(k)$ reaches the $k^{-5/3}$ scaling at a small range of $k$ around $600$ $h$Mpc-1. At lower redshifts, the $k$ range where this scaling holds increases, reaching as low as $\approx 200$ $h$Mpc-1 by $z=4$. In the bottom left, we show $E(k)$ at $z=5$ for simulations with varying spatial resolution. The scaling at turbulent scales is slightly shallower than $k^{-5/3}$ – the green dot-dashed curve is best fit by a $k^{-1.45}$ scaling. This may be because turbulent eddies do not fill the entire volume, that the scale of turbulent driving $L\approx c_{s}\Delta t$ is increasing with time, and/or that some non-turbulent kinetic motion is present. The fall-off in $E(k)$ shifts to the right with increasing resolution, confirming that this is due to numerical dissipation. In reality, when the turbulence is hydrodynamic the cascade should extend to $k_{\nu}\approx{\rm Re}^{3/4}/L$ where the Reynolds number for mean density gas is given by
×$\times\left(\frac{1+z}{7}\right)^{3}\left(\frac{T}{2\times 10^{4}{\rm K}}\right)^{-5/2},$
where $V_{\rm dr}$ is the velocity of the driving and $\nu$ is the kinematic viscosity for mean-density ionized gas. This indicates that the hydrodynamic turbulence can extend to sub-parsec scales for kiloparsec driving and applicable parameters. As discussed in the next section, the short turnover time of these smaller eddies (with the smallest eddies turning over a time $\text{Re}^{-1/2}$ times shorter than the largest) allows turbulence to grow magnetic fields even though we find that the smallest eddies we can resolve turnover only $N_{\rm\tau}\approx 1$ time by $z=4$. We find that $N_{\tau}$ is smaller than this when pre-heating by X-rays is significant (as low as $\approx 0.1$ for $T_{\min}=1000$ K, see Supplemental Material for more detail).
IV.2 The turbulent dynamo & B-field growth

### v2

@@ Abstract
The intergalactic medium (IGM) underwent intense heating that resulted in pressure disequilibrium in the wake of ionization fronts during cosmic reionization. The dynamical relaxation to restore pressure balance may have driven small-scale turbulence and, hence, the amplification of intergalactic magnetic fields. We investigate this possibility for the first time using a suite of $\approx 100$ pc resolution radiation-hydrodynamics simulations of IGM gas dynamics. We show that as the spatial resolution improves beyond that achieved with most prior studies, much of the IGM becomes turbulent unless it was pre-heated to $\gg 100$~K before reionization. In our most turbulent simulations, we find that the gas energy spectrum follows the expected $k^{-5/3}$ Kolmogorov scaling to the simulation’s resolution, and the eddy turnover time of the turbulence is <1<1 Gyr at $k\approx 1$~kpc-1. Turbulence will grow magnetic fields, and we show that the fields grown by reionization-driven turbulence could explain lower limits on the strength of volume-filling B-fields from observations of TeV blazars. As reionization sweeps over the cosmos, this mechanism could create turbulence throughout the cosmic volume with a character that only depends on the amount of IGM preheating.
@@ I Introduction
Hydrogen in the Intergalactic Medium (IGM) underwent a major phase transition about a billion years after the big bang, when it was ionized by the first generation of galaxies. This transition not only changed its ionization state, but also its temperature [1, 2, 3] and small-scale dynamics [4, 5, 6, 7, 8]. Before reionization, models predict that the IGM temperature was in the range 1010-10001000~[9, 10, 11]. For the coldest of these temperatures, the gas may have formed structures as light as $10^{4}~M_{\odot}$ [12, 13]. Reionization impulsively heated this gas to a few $\times 10^{4}$K, over-pressurizing these systems and causing them to dynamically relax to a scale of tens of kiloparsecs over a few hundred million years [6].
@@
Several studies have used high-resolution cosmological simulations to study this relaxation process and how it affects various observables  [5, 14, 6, 15, 16, 17, 18]. Here we investigate whether relaxation could drive turbulence that was unresolved in these prior studies. There is reason to suspect that the post-reionization IGM may have been turbulent. After ionization fronts (boundaries between fully neutral and highly ionized gas, I-fronts) sweep through a region, the gas would have been evacuated out of mini-halos and filaments and transported $\sim 10$s of co-moving kiloparsecs [6, 8], providing a natural driving mechanism. As the recently ionized gas expands out of the potential well, it cools adiabatically, reaching temperatures as low as a few thousand K [14, 17]. Nearby diffuse gas is compressed by this process, and can reach $5\times 10^{4}$~K or higher within $\approx 100$ Myr of reionization. The pressure differences between hot and cool gas can then drive turbulence. This potential mechanism has the unique features that it would be expected to occur everywhere in the IGM (since reionization happens everywhere) and be independent of process at play within galaxies and the circumgalactic medium. To our knowledge, a mechanism with these properties has never been studied before.
@@
Turbulent conditions in the IGM during and after reionization may grow intergalactic magnetic fields. There has been significant previous research on the seeding of IGM magnetic fields. One of the most referenced seeding mechanisms is the “Biermann battery”, in which inhomogeneous heating leads to small charge separation and the resulting electric fields generate B fields via Faraday’s law. This can occur during reionization [19, 20, 21] or in the pre-recombination epoch [22, 23]. Additionally, there may be seeds generated prior to reionization, such as from the streaming of cosmic rays produced by the first supernovae [24] and possibly even primordially during the Big Bang [25]. However, regardless of whether these seed fields are produced before recombination (“primordial” seed fields) or afterwards (“astrophysical” seed fields), their expected amplitudes are $\lesssim 10^{-20}$[e.g. 26, 23, 27], their amplitudes are expected to be small, $\lesssim 10^{-20}$ G. Reaching larger field strengths requires an amplification mechanism. One such mechanism is the so-called turbulent dynamo, which has been explored in detail both in general setups [28, 29, 30] and, e.g. in the context of forming proto-galaxies [31, 32, 29]. Here we consider the turbulent dynamo in the context of reionization-driven turbulence.
@@
Explaining the growth of volume-filling IGM magnetic fields is of particular interest in light of lower limits on the strength of these fields from TeV blazars [33, 34]. In the absence of magnetic fields, one expects to observe $\gamma$-ray halos around these objects generated by the inverse-Compton scattering of TeV photons off of CMB photons. The absence of these halos both observed in the spectral energy distribution and the spatial extent of the sources has been interpreted as an effect of deflection by IGM magnetic fields (an alternative plasma instability hypothesis has been recently ruled out  [^1] An alternative explanation of the blazar observations is that the TeV pairs lose their energy owing to plasma beam instabilities [62]. After an extensive debate, this possibility appears to have been ruled out. It was recently shown that Landau damping by non-thermal electrons generated by Compton scattering off of the gamma ray background should suppress the growth of these instabilities [63].). This absence places a conservative lower limit on their field strength of $B>3\times 10^{-16}$ G for fields correlated over $10$ kpc [36, 37, 38]. This explanation requires significant growth of seed fields, whatever their origin. It also requires that the mechanism responsible for driving turbulence be effective everywhere in the IGM to explain Blazar observations. Existing mechanisms for producing such turbulence [e.g. 39, 40] rely on structure formation and/or the galactic winds, and so can only produce IGM turbulence close to collapsed structures. We explore in this Letter whether this can be achieved by turbulence driven by reionization  [^2] Galactic and AGN feedback might be the other possibility [64], although this is not anticipated to be volume filling..
@@ II Numerical Methods
Our simulations of small-scale IGM gas dynamics were run using a modified version of the RadHydro code of Trac and Pen [43] and Trac and Cen [44]. Our version of RadHydro performs plane-parallel ray tracing radiative transfer (RT), with rays traced from the boundaries of cubical sub-volumes (RT domains) within our simulation box. Gravity is solved using the particle-mesh (PM) method. Our simulations are initialized at $z=300$ using separate transfer functions for dark matter and baryons from CAMB [45] and are run to $z=4$. Our fiducial box size is $L_{\rm box}=250$ $h^{-1}$kpc, with $1024^{3}$ gas cells, RT cells, and dark matter particles, for a fiducial spatial resolution of $\Delta x=244$ $h^{-1}$pc. Boundary conditions for the hydrodynamics and gravity are periodic  [^3] Note that because of the construction of the RT domains (in which rays terminate that edges of domains), explicitly periodic boundary conditions for the RT is not necessary. .
We run the simulation with hydrodynamics and gravity only (no RT) until the reionization redshift, $z_{\rm re}$. Plane-parallel ionization fronts (I-fonts) cross the RT domains at a redshift $z_{\rm re}$, following the procedure described in Cain et al. [17] (see also §C.2 in supplemental material). We use the reduced speed of light approximation to save computation time after I-fronts have crossed the RT domains. Following D’Aloisio et al. [6], the ionizing spectrum is discretized into $5$ frequency bins between $1$ and $4$ Ry, with a power law dependence $J_{\nu}\propto\nu^{-1.5}$. The strength of the ionizing background is set such that the HI photo-ionization rate is $\Gamma_{\rm HI}=3\times 10^{-13}$ s-1 everywhere (except within self-shielded systems). To study numerical convergence, we have also run a simulation with $2048^{3}$ grid cells and dark matter particles ($\Delta x=122$ $h^{-1}$pc), and ones with lower resolution  [^4] Due to computational limitations, we flash-ionized this $2048^{3}$ run instead of tracking RT (that is, we imposed a uniform ionizing background everywhere at $z=z_{\rm re}$). We show in the supplemental material that this approximation does not significantly alter the subsequent relaxation.
Figure 1: Snapshots showing the evolution of the density (top panels) and temperature (bottom panels) in our fiducial simulation. Immediately after ionization ($z=6.99$), the gas is clumpy at sub-kpc scales, reflecting its pre-ionized cold temperature, and is nearly isothermal, reflecting the heat input from reionization. Subsequently, over-dense filaments expand, generating a medium of expansion-cooled and compression-heated gas. By $z=6$, the gas begins to show signs of turbulence, and at $z\leq 5.5$, turbulence is observed throughout, in both the density and temperature. As relaxation slows, the turbulence dissipates, but some still persists to at least $z=4$. The inset in the bottom right shows the physical scale.
@@ III.1 Dynamics
Figure 1 shows the time evolution of the IGM gas density (top panels) and temperature (bottom panels) in our fiducial simulation, where reionization occurs at a redshift of $z_{\rm re}=7$. At $z=6.99$, the gas has had negligible time to respond to the photo-heating and retains its initial clumpiness. The low-density gas is nearly isothermal, at $\approx 30,000$~K, while the over-dense filaments are slightly cooler. These temperature differences result from the I-fronts moving more slowly through filaments, since the resulting temperature scales with the local I-front speed  [2, 3].
@@
Figure 2: Comparison of turbulence in runs with and without X-ray pre-heating, showing our highest resolution $0.25~h^{-1}$~Mpc, $2048^{3}$ simulations. Left: slice through the temperature field at $z=5$. The simulation assumes no heating before reionization such that the thermal evolution is essentially adiabatic. Right: the same, but for a run with a minimum gas temperature of $100$ K imposed at $z<15$. We see less turbulence in this case, with it being concentrated largely around the filaments.
@@
Between $z=7$ and $6.5$, the gas dynamics are consistent with the findings of Cain et al. [17] (see also Hirata [14]). The filamentary structures expand and cool adiabatically in response to impulsive heating from reionization. Meanwhile, the gas surrounding the expanding filaments is compressed and heated to $\gtrsim 50,000$~K. By $z=6.5$, there is a factor of $\approx 10$ temperature difference between over-dense, expanding gas and surrounding hot, compressed gas at lower densities. At $z=6$, these two phases are beginning to turbulently mix, and, by $z=5.5$, turbulence has developed, and is clearly seen in the density and temperature fields. Some of the turbulence may be instigated by Rayleigh-Taylor instabilities of less dense but over-pressurized gas expanding into denser but lower pressure regions. Regardless, it is clear that the driving mechanism of the turbulence in these simulations is the pressure-smoothing of filaments and mini-halos and the resulting overlap of gas evacuating these structures. By $z=4$, $700$ Myr after $z_{\rm re},$ relaxation is nearly complete, and the turbulence has started to dissipate.
@@
No previous works (to our knowledge) have observed turbulent behavior in the context of the reionized, diffuse IGM. We believe this is because they did not achieve the required spatial resolution. The highest-resolution studies to-date in this context (using grid or particle-based codes) achieve resolutions of $\approx$ 121-2 ckpc [5, 14, 6, 8, 17]. The focus of these studies was resolving the pre-ionization Jeans scale – the smallest scale at which the gas clumps –, which corresponds to $\approx 1-10$kpc (see Appendix A of [6]). We have run a convergence test (see supplemental material) and found that turbulent features are not seen at all in our simulations for cell sizes larger than $0.5$ $h^{-1}$kpc. We show in Appendix A that this resolution requirement is set not only by the Jeans scale of the ionized gas, but also by the scale at which numerical discretization introduces effective viscosity that prevents turbulence from forming [48]. We further show that in the context of post-reionization pressure smoothing, resolutions of 0.510.5-1 kpc are the expected minimum requirement to start resolving turbulence.
@@ III.2 Effect of X-ray pre-heating
Pre-heating by X-rays before reionization increases the Jeans scale of the pre-ionized gas above that of the limit of pure “adiabatic” evolution that is assumed by our fiducial simulation  [^5] Note that Compton heating off the CMB is significant at $z\gtrsim 100$, and we do include this effect. At $z\lesssim 100$, cosmic expansion and X-ray pre-heating at the only significant heating/cooling channels before reionization. . Smoothing of small-scale structure by pre-heating could inhibit the formation of turbulence by eliminating the smallest structures and making it less likely for expanding filaments to overlap. Upper limits on the 21 cm power spectrum from HERA [50] already constrain the temperature of the neutral IGM to be $\gtrsim 10$ K by $z\approx 8$. Theoretical expectations for the minimum temperature, $T_{\min}$, of the pre-heated IGM are as high as $T_{\min}\approx 1000$[e.g. 51]. However, recent work presenting more careful modeling of the spectra of high-redshift X-ray binaries (the most plausible source of such heating) anticipates temperatures closer to $\approx 100$ K near the start of reionization [9]. We model the effects of X-ray heating in an approximate manner, following D’Aloisio et al. [6], by imposing a temperature floor of $T_{\min}$ at $z<15$. Our pre-heating implementation likely overestimates the effect of pre-heating on structure formation for a given $T_{\min}$, since X-ray pre-heating is likely to trace the exponential growth of structure formation [e.g. 52].
@@
In Figure 2, we show slices through the temperature at $z=5$ for our highest-resolution adiabatic simulation with $122$ $h^{-1}$pc resolution (left), and the same run but with $T_{\min}=100$ K (right). As in Figure 1, we see turbulence throughout the box in the left panel. On the right, however, turbulent features are much more sparse, favoring regions around the most massive filaments. The sparsity of dense structures before reionization makes it less likely once reionized for their photo-evaporating remnants to collide, which suppresses the driving mechanism of the turbulence. We note that the formation of massive structures, around which turbulence still forms in the $T_{\min}=100$K run, is suppressed in our tiny boxes. As such, the fraction of gas that is turbulent in our $T_{\min}=100$~K run is likely an underestimate for the entire IGM. Appendix B and the supplementary material (§C.1.4) presents a $T_{\min}=1000$~K simulation, in which there is no evidence for turbulence.
IV The turbulent dynamo & B-field growth
Turbulence will act to grow magnetic fields via the turbulent dynamo. We show in Appendix B that our simulations exhibit the Kolmogorov energy spectrum scaling expected for turbulent gas ($\propto k^{-5/3}$), down to a maximum $k$ set by the resolution of the simulation (see Appendix A). In the absence of this “numerical viscosity”, we expect to see turbulence extend down to a minimum scale set by the viscosity of intergalactic hydrogen, which works out to a co-moving scale of about a parsec for our estimated Reynolds number of ${\rm Re}\sim 10^{5}$ (Appendix B).